if a 12' flagpole casts an 18' shadow at the same time that a nearby building casts a 72' shadow, how tall is the building?
Answer by checkley71(8403)

12/18=X/72 OR 18X=12*72 OR 18X=900 OR X=800/18 OR X=50'